# Lifting subgroups of $\mathrm{PSL}_{2}$ to $\mathrm{SL}_{2}$ over local fields

### v1

@@ 1. Introduction
Here we consider the case of $\pi\colon{\rm SL_{2}}(K)\twoheadrightarrow{\rm PSL_{2}}(K)$ over a field $K$. If ${\rm char}(K)=2$, then ${\rm SL_{2}}(F)\cong{\rm PSL_{2}}(K)$, so every subgroup of ${\rm PSL_{2}}(K)$ lifts to ${\rm SL_{2}}(K)$. Hence we will assume ${\rm char}(K)\neq 2$. In this case, $-I$ is the unique element of order 2 in ${\rm SL_{2}}(K)$, so any subgroup of ${\rm PSL_{2}}(K)$ containing elements of order 2 necessarily cannot lift to ${\rm SL_{2}}(K)$.
Culler proved in [5, Theorem 3.1] that discrete subgroups of ${\rm PSL_{2}}(\mathbb{C})$ with no 2-torsion can be lifted to ${\rm SL_{2}}(\mathbb{C})$. Here we prove the following non-archimedean analogon of this result.
@@
We additionally obtain the following result in characteristic zero.
@@ 2. Background
Let $K$ be a non-archimedean local field with finite residue field of order $q=p^{r}$. We denote the valuation ring of $K$ by $\mathcal{O}_{K}$, the corresponding valuation map by $v$, and we let $u$ be a uniformiser of $K$. We use $T_{K}$ to denote the corresponding Bruhat–Tits tree, upon which ${\rm PSL_{2}}(K)$ (and hence ${\rm SL_{2}}(K)$) acts by isometries. Each element of ${\rm PSL_{2}}(K)$ (and their representatives in ${\rm SL_{2}}(K)$) can therefore be classified as either elliptic or hyperbolic, depending on its translation length on $T_{K}$. We will use ${\rm Fix}(g)$ to denote the fixed point set of an elliptic element $g$, and ${\rm Ax}(h)$ to denote the translation axis of a hyperbolic element $h$. We respectively refer to [4] and [10] for more information on non-archimedean local fields and the Bruhat–Tits tree.
@@ 2. $(2)$
Using the classification of finite subgroups of ${\rm PSL_{2}}(K)$ given in [11, Theorem 1], we obtain:
@@
Following [12, Section 2.4], we say that a group satisfying the second condition is of Borel type, as it can be represented in ${\rm SL_{2}}(K)$ by a group of upper triangular matrices. Since the boundary of $T_{K}$ can be identified with the projective line $\mathbb{P}^{1}(K)$ [10, p. 72], it follows that a group of Borel type necessarily fixes an end of $T_{K}$.
@@
As in [1], we define a band of radius $r$ and nerve $S$ to be the subtree of $T_{K}$ consisting of all vertices at distance at most $r$ from some geodesic $S$. Given a ray $R$ from a vertex $x$ to some end of $T_{K}$, we define a horoball to be the subtree of $T_{K}$ consisting of all vertices at distance at most $d(v,x)$ from each vertex $v$ of $R$.
@@
  • ${\rm Fix}(g)$ is a horoball, $\mathrm{char}(K)=p=n$ and ${\rm Fix}(g)={\rm Fix}(g^{i})$ for each $i\in\{1,\dots,p-1\}$.
@@
Let $\overline{g}$ be a representative of $g$ in ${\rm SL_{2}}(K)$, and note that ${\rm Fix}(g)={\rm Fix}(\overline{g})$.
If $\overline{g}$ is unipotent, then we may conjugate it into the form
@@
and ${\rm Fix}(g)$ is the horoball defined by the ray $\mathcal{O}_{K}e_{1}+u^{-v(x)+n}\mathcal{O}_{K}e_{2}$ for $n\geq 0$ [1, Proposition 20]. Since $g$ has finite order, $K$ must have characteristic $p$, whence $g$ has order $p$ and it is straightforward to check that each non-trivial power $g^{i}$ has the same horoball as its fixed point set.
Hence we may assume $\overline{g}$ is diagonalisable over either $K$ or a quadratic extension $K^{\prime}$ of $K$. In the first case, ${\rm Fix}(g)$ is a band with nerve the unique bi-infinite ray between the ends of $T_{K}$ determined by the eigenvectors of $\overline{g}$. In the latter case, we consider the bi-infinite ray $S^{\prime}$ between the ends of the Bruhat–Tits tree $T_{K^{\prime}}$ corresponding to $K^{\prime}$ determined by the eigenvectors of $\overline{g}$. The intersection $S^{\prime}\cap T_{K}$ is either a vertex or the midpoint of an edge, and that vertex/edge is the nerve of ${\rm Fix}(g)$; see [1, Proposition 18] for further details. ∎
@@ 2.2. Bass–Serre theory
  • For each vertex $v\in V(Y)$, a group $G_{v}$. If $T$ is a spanning tree of $Y$, then $G$ is isomorphic to the fundamental group $\pi_{1}(\mathbb{G},T)$. Also note that $G_{v}\cong G_{w}$ for all $w\in V(X)$ representing $v\in V(Y)$; see [10, I.5] for details. This is the situation we consider below.
@@
Theorem 2.7 ([10, Section 5.4, Theorem 13]).
@@ 3. Proofs of Theorems A and B
Let $G_{x}$ be the stabiliser in $G$ of either a vertex or an edge of $T_{K}$. By 2.1, $G_{x}$ is a finite group. First we show that at most one lift of $G_{x}$ to ${\rm SL_{2}}(K)$ exists. Let $g\in G_{x}$ have order $k$. Since $G$ and hence $G_{x}$ has no 2-torsion, $k$ is odd and exactly one of the pre-images $\pi^{-1}(g)=\{g^{\prime},-g^{\prime}\}$ in ${\rm SL_{2}}(K)$ has order $k$. Thus each element of $G_{x}$ has a unique lift, so a lift of $G_{x}$, if it exists, must be unique. Next we show that such a lift exists. By 2.2, $G_{x}$ is either cyclic or of Borel type. If $G_{x}$ is cyclic, then the lift of $G_{x}$ is determined by the lift of a generator of $G_{x}$. Hence we may suppose that $G_{x}$ is of Borel type and $\mathrm{char}(K)=p$. By conjugating if necessary, we may assume that $G_{x}$ fixes $\infty$ on the projective line $\mathbb{P}^{1}(K)$.
@@
We need to check that (for every edge $e$) $f_{1}=f_{2}$. Notice that the relation $t_{e}\sigma_{e}(g)t_{e}^{-1}=\sigma_{\overline{e}}(g)$ implies that $f_{2}(g)=\phi(t_{e}\sigma_{e}(g)t_{e}^{-1})$ for $g\in G_{e}$. Let $\theta$ be the inner automorphism of $G$ given by $\theta(h)=t_{e}^{-1}ht_{e}$ for each $h\in G$. Note that $\theta$ is surjective and it sends $\sigma_{\overline{e}}(G_{e})$ isomorphically to $\sigma_{e}(G_{e})$. Now consider the compositions $\pi\circ f_{1}\circ\theta$ and $\pi\circ f_{2}\circ\theta$, where we restrict to $G_{e}=\sigma_{\overline{e}}(G_{e})$ so that these maps are well-defined. Observe that, since $\pi$ is a homomorphism and $\pi\circ\phi$ is the identity map on $W$, both compositions are the identity map; in particular $f_{1}\circ\theta$ and $f_{2}\circ\theta$ are both lifts of $G_{e}$. Hence by 3.1 these two lifts are equal. Since $\theta$ is surjective, $f_{1}=f_{2}$ as we needed to show. ∎
@@
In general it is not true that a lift on vertex stabilisers can be extended to a lift of the whole group; the uniqueness appears to be necessary unless one has very fine control. Instructive counterexamples can be found in the quotient map from the Baumslag–Solitar group $BS(4,6)\to BS(2,3)$, as well as by considering the quotient of $\langle a:a^{2}\rangle\times\langle x,y:x^{3}=y^{3}\rangle$ to its right hand factor.
@@
Let $G$ be a discrete subgroup of ${\rm PSL_{2}}(K)$ with no 2-torsion, where $\mathrm{char}(K)=0$. Let $X$ be the tree obtained from $T_{K}$ by collapsing ${\rm Fix}(g)$ to a single vertex for each elliptic element $g\in G$. Note that the collapsing map $f\colon T_{K}\to X$ is equivariant, and so $X$ inherits an action of $G$ from $T_{K}$. We also observe that, since all fixed point sets were collapsed, every edge stabiliser of $X$ in $G$ is trivial. Now let $v$ be a vertex of $X$. We first show that $f^{-1}(v)={\rm Fix}(g)$ for some $g\in G$. Indeed, if $x,y\in f^{-1}(v)$, then there is some sequence of vertices $x=x_{0},x_{1},\dots,x_{n}=y$ of $T_{K}$ where each consecutive pair $\{x_{i},x_{i+1}\}$ is contained in some ${\rm Fix}(g_{i})$, and we may assume that $g_{i}\neq g_{i+1}$ (if not, $x_{i+1}$ may be omitted). This implies that $x_{i+1}\in{\rm Fix}(g_{i})\cap{\rm Fix}(g_{i+1})$, whence 2.6 implies that either ${\rm Fix}(g_{i})\subseteq{\rm Fix}(g_{i+1})$, or ${\rm Fix}(g_{i+1})\subseteq{\rm Fix}(g_{i})$. In either case, both $x_{i}$ and $x_{i+1}$ are contained in the larger fixed point set, and we may remove $x_{i+1}$ from the sequence. Continuing in this fashion, the sequence is eventually reduced to the two vertices $x$ and $y$, which must be contained in ${\rm Fix}(g)$ for some $g\in G$. Note that $g$ necessarily has finite order by 2.1.
Now let $H$ be the set-wise stabiliser in $G$ of $f^{-1}(v)={\rm Fix}(g)\subseteq T_{K}$. By 2.4, ${\rm Fix}(g)$ is a band with nerve $S$ which is either a vertex, an edge or a bi-infinite ray. Note that $H$ must preserve $S$. If $S$ is either a vertex or an edge, then $H$ is a finite cyclic group by 2.3. On the other hand, if $S$ is a bi-infinite ray, then each element of $H$ either fixes or swaps the ends of $S$. Since $G$ contains no elements of order 2, we deduce that every element of $H$ fixes both ends of $S$ and is hence either finite cyclic or a direct product of a finite cyclic and an infinite cyclic group, depending on whether or not there is a hyperbolic element in $G$ which fixes both ends of $S$. The conclusion follows by applying Theorem 2.7: since edge stabilisers are trivial, the edge relations in the fundamental group are empty, and so $G$ is a free product of its vertex stabilisers and a free group. ∎
@@
The difficulty with duplicating this argument in positive characteristic lies in the subgroups of Borel type, and understanding how their fixed point sets can intersect fixed point sets of other subgroups.
@@ 4. Proof of Theorem C
We recall the classication of local fields [9, Proposition II.5.2]. Let $K$ be a non-archimedean local field. If $\mathrm{char}(K)=0$, then $K$ is a finite extension of $\mathbb{Q}_{p}$ and $\mathbb{Q}$ is the corresponding prime field. If $\mathrm{char}(K)>0$, then $K$ is a field of formal Laurent series $\mathbb{F}_{q}((t))$ and $\mathbb{F}_{p}$ is the corresponding prime field. The following two basic lemmas are well-known.
@@
Observe that $A$ and $B$ generate $Q_{8}$, the quaternion group with order $8$, and that $ACA^{-1}=BCB^{-1}=C^{-1}$. Moreover, [8, Proposition II.3.15] implies that $C$ is hyperbolic (with respect to $T_{K(i)}$) and hence of infinite order. The respective images $\alpha,\beta,\gamma,\delta$ of $A,B,C,D$ in ${\rm PSL_{2}}(K(i))$ generate a subgroup with the following presentation:
@@
A similar argument to the above rules out the parities where $c=d=0$ or $b=c=d=1$. In the second set of matrices, the roles of $A$ and $B$ are interchanged, while retaining the property that $A$ is diagonal. Observe that $A$ is hyperbolic by [8, Proposition II.3.15], and that $CAC^{-1}=DAD^{-1}=A^{-1}$. Hence similar reasoning shows that this rules out the parities $(a,b,c,d)=(1,0,1,1)$.

### v2

@@ 1. Introduction
Here we consider the case of $\pi\colon{\rm SL_{2}}(K)\twoheadrightarrow{\rm PSL_{2}}(K)$ over a field $K$. If ${\rm char}(K)=2$, then ${\rm SL_{2}}(K)\cong{\rm PSL_{2}}(K)$, so every subgroup of ${\rm PSL_{2}}(K)$ lifts to ${\rm SL_{2}}(K)$. Hence we will assume ${\rm char}(K)\neq 2$. In this case, $-I$ is the unique element of order 2 in ${\rm SL_{2}}(K)$, so any subgroup of ${\rm PSL_{2}}(K)$ containing elements of order 2 necessarily cannot lift to ${\rm SL_{2}}(K)$.
Let $K$ be a finitely generated subgroup of $\mathrm{SL}(2,\bar{\mathbb{Q}})$, where $\bar{\mathbb{Q}}$ are the algebraic numbers. Even though $K$ might not be discrete in $\mathrm{SL}(2,\mathbb{C})$ it embeds discretely in a finite direct product with factors $\mathrm{SL}(2,k_{i})$ where the $k_{i}$ are local fields determined by the finitely many entries of the generators.
Each local field is either archimedean (and isomorphic to $\mathbb{R}$ and $\mathbb{C}$ with their standard topologies) or non-archimedean (and totally disconnected).
Culler proved in [5, Theorem 3.1] that discrete subgroups of ${\rm PSL_{2}}(\mathbb{C})$ with no 2-torsion can be lifted to ${\rm SL_{2}}(\mathbb{C})$. Here we prove the following non-archimedean analogon of this result. Although every non-archimedean local field of characteristic zero (i.e. a finite extension of $\mathbb{Q}_{p}$) has the same first order theory as $\mathbb{C}$, the differing topologies mean that, even in characteristic 00, we require a different approach here than that used by Culler.
@@
We additionally obtain an analogon of Fricke’s result that many Fuchsian groups, including all non-cocompact ones, are free products of a free group and a free product of cyclic groups amalgamated over an infinite cyclic subgroup [9, p. 98]. For Kleinian groups the situation is far more complicated.
@@
Remark 1.1.
Throughout the paper we refer to elements of ${\rm PSL_{2}}(K)$ as unipotent if one of their lifts is; in particular, if the lifts are $\{U,-U\}$ for some unipotent element $U$ of ${\rm SL_{2}}(K)$.
@@ 2. Background
Let $K$ be a non-archimedean local field with finite residue field of order $q=p^{r}$. We denote the valuation ring of $K$ by $\mathcal{O}_{K}$, the corresponding valuation map by $v$, and we let $u$ be a uniformiser of $K$. We use $T_{K}$ to denote the corresponding Bruhat–Tits tree, upon which ${\rm PSL_{2}}(K)$ (and hence ${\rm SL_{2}}(K)$) acts by isometries, without inversion. Each element of ${\rm PSL_{2}}(K)$ (and their representatives in ${\rm SL_{2}}(K)$) can therefore be classified as either elliptic or hyperbolic, depending on its translation length on $T_{K}$. We will use ${\rm Fix}(g)$ to denote the fixed point set of an elliptic element $g$, and ${\rm Ax}(h)$ to denote the translation axis of a hyperbolic element $h$. We respectively refer to [4] and [12] for more information on non-archimedean local fields and the Bruhat–Tits tree.
@@ 2. $(2)$
Using the classification of finite subgroups of ${\rm PSL_{2}}(K)$ given in [13, Theorem 1], we obtain:
@@
Following [14, Section 2.4], we say that a group satisfying the second condition is of Borel type, as it can be represented in ${\rm SL_{2}}(K)$ by a group of upper triangular matrices. Since the boundary of $T_{K}$ can be identified with the projective line $\mathbb{P}^{1}(K)$ [12, p. 72], it follows that a group of Borel type necessarily fixes an end of $T_{K}$.
@@
As in [1], we define a band of radius $r$ and nerve $S$ to be the subtree of $T_{K}$ consisting of all vertices at distance at most $r$ from some geodesic $S$. Given an infinite ray $R$ emanating from a vertex $x$ of $T_{K}$, we define a horoball to be the subtree of $T_{K}$ consisting of all vertices at distance at most $d(v,x)$ from each vertex $v$ of $R$.
@@
  • ${\rm Fix}(g)$ is a horoball and $\mathrm{char}(K)=p=n$.
@@
Let $\overline{g}$ be a representative of $g$ in ${\rm SL_{2}}(K)$, and note that ${\rm Fix}(g)={\rm Fix}(\overline{g})$. If $\overline{g}$ is unipotent, then we may conjugate it into the form
@@
and ${\rm Fix}(g)$ is the horoball defined by the ray $\mathcal{O}_{K}e_{1}+u^{-v(x)+n}\mathcal{O}_{K}e_{2}$ for $n\geq 0$ [1, Proposition 20]. Since $g$ has finite order, $K$ must have characteristic $p$, whence $g$ has order $p$. We may assume $\overline{g}$ is diagonalisable over either $K$ or a quadratic extension $K^{\prime}$ of $K$. In the first case, ${\rm Fix}(g)$ is a band with nerve the unique bi-infinite ray between the ends of $T_{K}$ determined by the eigenvectors of $\overline{g}$. In the latter case, we consider the bi-infinite ray $S^{\prime}$ between the ends of the Bruhat–Tits tree $T_{K^{\prime}}$ corresponding to $K^{\prime}$ determined by the eigenvectors of $\overline{g}$. The intersection $S^{\prime}\cap T_{K}$ is either a vertex or the midpoint of an edge, and that vertex/edge is the nerve of ${\rm Fix}(g)$; see [1, Proposition 18] for further details. ∎
@@ 2.2. Bass–Serre theory
  • For each vertex $v\in V(Y)$, a group $G_{v}$. If $T$ is a spanning tree of $Y$, then $G$ is isomorphic to the fundamental group $\pi_{1}(\mathbb{G},T)$. Also note that $G_{v}\cong G_{w}$ for all $w\in V(X)$ representing $v\in V(Y)$; see [12, I.5] for details. This is the situation we consider below.
@@
Theorem 2.7 ([12, Section 5.4, Theorem 13]).
@@
Collapsing Maps
Given an action of $G$ on a tree $T$, we can obtain a new action of $G$ on another tree $T^{\prime}$ by collapsing each member of some $G$-orbit(s) of edges to a point. This process usually changes the elliptic subgroups: the effect on the graph of groups is to replace any collapsed subgraphs with their (Bass–Serre theoretic) fundamental group, which up to isomorphism is the setwise stabiliser of some largest subtree containing only edges which are collapsed. If one has sufficient control over these subtrees and their setwise stabilisers, then this can clarify the algebraic structure of $G$. There is a $G$-equivariant collapse map $f:T\to T^{\prime}$ given by sending each edge and vertex in a collapsed subtree to the vertex corresponding to it. (See for instance [6, Section 3] for a discussion of these and related ideas involved in comparing actions of the same group on different trees.)
@@ 3. Proofs of Theorems A and B
Let $G_{x}$ be the stabiliser in $G$ of either a vertex or an edge of $T_{K}$. By 2.1, $G_{x}$ is a finite group. First we show that at most one lift of $G_{x}$ to ${\rm SL_{2}}(K)$ exists. Let $g\in G_{x}$ have order $k$. Since $G$ and hence $G_{x}$ has no 2-torsion, $k$ is odd and exactly one of the pre-images $\pi^{-1}(g)=\{g^{\prime},-g^{\prime}\}$ in ${\rm SL_{2}}(K)$ has order $k$. Thus each element of $G_{x}$ has a unique lift, so a lift of $G_{x}$, if it exists, must be unique.
Next we show that such a lift exists. By 2.2, $G_{x}$ is either cyclic or of Borel type. If $G_{x}$ is cyclic, then the lift of $G_{x}$ is determined by the lift of a generator of $G_{x}$. Hence we may suppose that $G_{x}$ is of Borel type and $\mathrm{char}(K)=p$. By conjugating if necessary, we may assume that $G_{x}$ fixes $\infty$ on the projective line $\mathbb{P}^{1}(K)$.
@@
We need to check that (for every edge $e$) $f_{1}=f_{2}$. Notice that the relation $t_{e}\sigma_{e}(g)t_{e}^{-1}=\sigma_{\overline{e}}(g)$ implies that $f_{2}(\sigma_{e}(g))=\phi(t_{e}\sigma_{e}(g)t_{e}^{-1})$ for $g\in G_{e}$. Let $\theta$ be the inner automorphism of $G$ given by $\theta(h)=t_{e}^{-1}ht_{e}$ for each $h\in G$. Note that $\theta$ is surjective and it sends $\sigma_{\overline{e}}(G_{e})$ isomorphically to $\sigma_{e}(G_{e})$. Now consider the compositions $\pi\circ f_{1}\circ\theta$ and $\pi\circ f_{2}\circ\theta$, where we restrict to $G_{e}=\sigma_{\overline{e}}(G_{e})$ so that these maps are well-defined. Observe that, since $\pi$ is a homomorphism and $\pi\circ\phi$ is the identity map on $W$, both compositions are the identity map; in particular $f_{1}\circ\theta$ and $f_{2}\circ\theta$ are both lifts of $G_{e}$. Hence by 3.1 these two lifts are equal. Since $\theta$ is surjective, $f_{1}=f_{2}$ as we needed to show. ∎
@@
In general it is not true that a lift on vertex stabilisers can be extended to a lift of the whole group; the uniqueness appears to be necessary unless one has very fine control. Consider the quotient map from the Baumslag–Solitar group $\mathrm{BS}(4,6)=\langle a,t:ta^{4}t^{-1}=a^{6}\rangle$ to $\mathrm{BS}(2,3)=\langle b,s:sb^{2}s^{-1}=b^{3}\rangle$, sending $a\mapsto b$ and $t\mapsto s$. Note that $\mathrm{BS}(2,3)$ acts on a tree with one orbit each of edges and vertices (as an HNN extension of $\mathbb{Z}$ with respect to its index 2 and 3 subgroups), and both the vertex group and edge generators lift to the generators of $\mathrm{BS}(4,6)$, but this cannot extend to a lift of the whole group since the groups are not isomorphic. The failure of uniqueness is witnessed by elements such as $ta^{2}t^{-1}a^{-2}$, which is a loxodromic element of $\mathrm{BS}(4,6)$ descending to $b$ in $\mathrm{BS}(2,3)$.
@@
Let $G$ be a discrete subgroup of ${\rm PSL_{2}}(K)$ with no 2-torsion, where $\mathrm{char}(K)=0$. Let $X$ be the tree obtained from $T_{K}$ by collapsing every edge with non-trivial stabiliser. Note that the collapsing map $f\colon T_{K}\to X$ is equivariant, and so $X$ inherits an action of $G$ from $T_{K}$.
We observe that, since all fixed point sets were collapsed, every edge stabiliser of $X$ in $G$ is trivial.
Now let $v$ be a vertex of $X$. We first show that $f^{-1}(v)={\rm Fix}(g)$ for some $g\in G$. Indeed, if $x,y\in f^{-1}(v)$, then there is a path joining them where every edge has non-trivial stabiliser. In particular, there is some sequence of vertices $x=x_{0},x_{1},\dots,x_{n}=y$ of $T_{K}$ where each consecutive pair $\{x_{i},x_{i+1}\}$ is contained in some ${\rm Fix}(g_{i})$, and we may assume that $g_{i}\neq g_{i+1}$ (if not, $x_{i+1}$ may be omitted). This implies that $x_{i+1}\in{\rm Fix}(g_{i})\cap{\rm Fix}(g_{i+1})$, whence 2.6 implies that either ${\rm Fix}(g_{i})\subseteq{\rm Fix}(g_{i+1})$, or ${\rm Fix}(g_{i+1})\subseteq{\rm Fix}(g_{i})$. In either case, both $x_{i}$ and $x_{i+1}$ are contained in the larger fixed point set, and we may remove $x_{i+1}$ from the sequence. Continuing in this fashion, the sequence is eventually reduced to the two vertices $x$ and $y$, which must be contained in ${\rm Fix}(g)$ for some $g\in G$. Note that $g$ necessarily has finite order by 2.1.
Now let $H$ be the set-wise stabiliser in $G$ of $f^{-1}(v)={\rm Fix}(g)\subseteq T_{K}$. By 2.4, ${\rm Fix}(g)$ is a band with nerve $S$ which is either a vertex, an edge or a bi-infinite ray. Note that $H$ must preserve $S$. If $S$ is either a vertex or an edge, then $H$ is a finite cyclic group by 2.3. On the other hand, if $S$ is a bi-infinite ray, then each element of $H$ either fixes or swaps the ends of $S$. Since $G$ contains no elements of order 2, we deduce that every element of $H$ fixes both ends of $S$. Hence we obtain a set of diagonalisable matrices with the same eigenvectors, so they commute and are simultaneously diagonalisable. If $H$ contains no hyperbolic elements then $H$ fixes a point of $S$ and is hence cyclic by 2.3. If $H$ does contain hyperbolic elements stabilising $S$ then it contains one of smallest translation length, hence the hyperbolic elements of $H$ form an infinite cyclic group. Since $H$ is abelian it is hence either a finite cyclic group or a direct product of a finite cyclic and an infinite cyclic group. The conclusion follows by applying Theorem 2.7: as remarked above the proof edge stabilisers are trivial, hence the edge relations in the fundamental group are empty, and so $G$ is a free product of its vertex stabilisers and a free group. ∎
@@
The difficulty with duplicating this argument in positive characteristic lies in the subgroups of Borel type, and understanding how their fixed point sets can intersect fixed point sets of other subgroups: it is no longer necessarily true that the subtrees $f^{-1}(v)$ coincide with fixed point sets.
@@ 4. Proof of Theorem C
We recall the classification of local fields [11, Proposition II.5.2]. Let $K$ be a non-archimedean local field. If $\mathrm{char}(K)=0$, then $K$ is a finite extension of $\mathbb{Q}_{p}$ and $\mathbb{Q}$ is the corresponding prime field. If $\mathrm{char}(K)>0$, then $K$ is a field of formal Laurent series $\mathbb{F}_{q}((t))$ and $\mathbb{F}_{p}$ is the corresponding prime field. The following two basic lemmas are well-known.
@@
Observe that $A$ and $B$ generate $Q_{8}$, the quaternion group with order $8$, and that $ACA^{-1}=BCB^{-1}=C^{-1}$. Moreover, [10, Proposition II.3.15] implies that $C$ is hyperbolic (with respect to $T_{K(i)}$) and hence of infinite order. The respective images $\alpha,\beta,\gamma,\delta$ of $A,B,C,D$ in ${\rm PSL_{2}}(K(i))$ generate a subgroup with the following presentation:
@@
A similar argument to the above rules out the parities where $c=d=0$ or $b=c=d=1$. In the second set of matrices, the roles of $A$ and $B$ are interchanged, while retaining the property that $A$ is diagonal. Observe that $A$ is hyperbolic by [10, Proposition II.3.15], and that $CAC^{-1}=DAD^{-1}=A^{-1}$. Hence similar reasoning shows that this rules out the parities $(a,b,c,d)=(1,0,1,1)$.